# Look Before You Leap: Enhancing Attention and Vigilance Regarding Harmful Content with GuidelineLLM

### v1

@@ Abstract
Despite being empowered with alignment mechanisms, large language models (LLMs) are increasingly vulnerable to emerging jailbreak attacks that can compromise their alignment mechanisms. This vulnerability poses significant risks to the real-world applications. Existing work faces challenges in both training efficiency and generalization capabilities (i.e., Reinforcement Learning from Human Feedback and Red-Teaming). Developing effective strategies to enable LLMs to resist continuously evolving jailbreak attempts represents a significant challenge. To address this challenge, we propose a novel defensive paradigm called GuidelineLLM, which assists LLMs in recognizing queries that may have harmful content. Before LLMs respond to a query, GuidelineLLM first identifies potential risks associated with the query, summarizes these risks into guideline suggestions, and then feeds these guidelines to the responding LLMs. Importantly, our approach eliminates the necessity for additional safety fine-tuning of the LLMs themselves; only the GuidelineLLM requires fine-tuning. This characteristic enhances the general applicability of GuidelineLLM across various LLMs. Experimental results demonstrate that GuidelineLLM can significantly reduce the attack success rate (ASR) against the LLMs (an average reduction of 34.17% ASR) while maintaining the helpfulness of the LLMs in handling benign queries. Code is available at https://github.com/sqzhang-lazy/GuidelineLLM.
@@ Introduction
To enhance the safety of LLMs, common approaches involve reinforcing their alignment mechanisms. Techniques such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022a) and Red-Teaming (Ganguli et al. 2022; Perez et al. 2022; Touvron et al. 2023) have proven effective in improving the ability of LLMs to detect harmful content. However, these methods demand extensive training data and significant computational resources, which can limit the speed of iterative improvements. Furthermore, additional retraining might degrade the performance of established safety alignment mechanisms (Qi et al. 2023; Zhou et al. 2024). Therefore, it is critical to investigate methodologies that can mitigate the production of harmful content without requiring modifications to the model parameters.
@@
Two prevalent methods are currently employed to enhance the safety policies of LLMs: alignment and red teaming. The objective of alignment is to bridge the gap between LLMs’ language modeling goals, such as predicting the next token during pre-training, and their ultimate aim of “following instructions and being helpful, truthful, and harmless” in practical applications (Ouyang et al. 2022). Common techniques in this domain include Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022a), Constitutional AI (Bai et al. 2022b), and self-alignment (Sun et al. 2024). These approaches focus on embedding alignment rules within pre-trained models to restrict harmful behaviors during inference. In LLMs research, the term “Red-Teaming” has recently come to denote the systematic testing or probing of LLMs to identify their potential for harmful behavior and uncover safety vulnerabilities. The critical aspect of red teaming is the aggregation of sufficient and diverse attacks, enabling LLMs to refuse responses to these attacks. Current efforts in this area mainly concentrate on automating the generation of high-quality, provocative queries, allowing LLMs to encounter a wider variety and richer range of attacks, thereby bolstering their defensive capabilities (Deng et al. 2023a; Ge et al. 2023; Jiang et al. 2024). However, recent studies also demonstrate that retraining the model can lead to the deterioration of safety alignment mechanisms (Qi et al. 2023; Zhou et al. 2024). This indicates that arbitrary parameter updates to the model are not advisable and can slow the model iteration process.
@@
Ensuring that LLMs output only safe content during inference is critically important. Common methodologies to achieve this include preprocessing the input before inference, employing specialized inference paradigms, and detecting harmful content through input-output result analysis. Given the sensitivity of LLMs to input prompts, modifying the prompt content can enhance their ability to focus on harmful content. For instance, incorporating a system prompt can assist LLMs in identifying harmful content (Xie et al. 2023). By leveraging LLMs’ in-context learning capabilities, models can be conditioned to recognize and refuse harmful content (Wei, Wang, and Wang 2023). Additionally, creating multiple perturbed samples of the input and aggregating the results has proven effective (Robey et al. 2023; Cao et al. 2023). Some approaches involve having LLMs analyze the intent of the question before providing an answer (Zhang et al. 2024; Zeng et al. 2024b). Alternatively, decoding strategies can be applied to suppress harmful prompts during generation, thereby ensuring safer outputs (Zhong et al. 2024). In terms of detection and classification, Meta has specifically fine-tuned a model named Llama-Guard, which can classify inputs and input-output pairs according to taxonomy and risk guidelines (Inan et al. 2023). In this work, we primarily focus on safeguarding LLMs during the inference phase. We aim to find a defensive paradigm that applies to different LLMs and can be continuously expanded to counter the ever-evolving jailbreak attacks.
@@
The prompts used to fine-tune GuidelineLLM consist of jailbreak and benign queries as inputs, with the corresponding guidelines as outputs. The prompts used to fine-tune JailbreakLLM require JailbreakLLM to generate attack queries and produce jailbreak queries based on specific jailbreak techniques. To increase data diversity, we sample two examples of the same jailbreak technique as inputs during the fine-tuning processes. During the inference phase, the prompts used by both GuidelineLLM and JailbreakLLM are consistent with those used in the fine-tuning phase.
@@
For harmful datasets, we select three in-the-wild jailbreak datasets (i.e., DAN (Shen et al. 2023), SAP200 (Deng et al. 2023a), JailbreakChatGPT (Liu et al. 2023b)) and a popular optimization-based automatic jailbreak method (i.e., AutoDAN (Liu et al. 2023a)). Additionally, we use gpt-3.5-turbo-0125 to generate a T-jailbreak dataset comprising 1,137 queries. Besides, to evaluate the impact of GuidelineLLM on the helpfulness of the responding LLMs, we sample 400 queries from the Alpaca and GPTTeacher datasets (Taori et al. 2023).
@@
Baseline
We select Llama2-7B-Chat (Touvron et al. 2023), Vicuna-7B, and Vicuna-13B (Chiang et al. 2023) as our responding LLMs. We compare the performance of our GuidelineLLM method with several other approaches: No defense, Self-Reminderr (Xie et al. 2023), IA (Zhang et al. 2024), and Llama3-Guard (Inan et al. 2023). For the Llama3-Guard, we only evaluate the harmfulness of the inputs. Considering that IA’s defensive method requires two rounds of inference, and both Self-Reminder and GuidelineLLM add additional prompts to the original query (i.e., Self-Reminder adds a system prompt, and GuidelineLLM adds guidelines), in the AutoDAN experiment, we use the input content from the last round of inference for each defensive method as the attack target.
@@
Llama2- 7B-Chat DAN - - - - - -
AutoDAN - - - - - -
T-Jailbreak 0 1.73 5.17 0 93.1 0.16
@@
Model Method A B C D E FRR Score
Llama2- 7B-Chat Vanilla 91.62 3.14 3.4 0.26 1.57 1.83 3.82
Guideline LLM 71.02 5.74 15.93 0.78 6.53 7.31 3.33
Vicuna- 7B Vanilla 91.27 0.79 5.03 0.26 2.65 2.91 3.77
Guideline LLM 75.94 5.35 14.71 0.53 3.48 4.01 3.50
Vicuna- 13B Vanilla 93.62 1.33 3.72 0.53 0.8 1.33 3.86
Guideline LLM 85.86 2.62 9.69 0.26 1.57 1.83 3.70
Table 4: The results of the helpfulness scores for GuidelineLLM and Vanilla. The results for A, B, C, D, E, and FRR are shown as percentages (%).
Model Benign Guideline A B C D E FRR Score
Llama2- 7B-Chat $\checkmark$ 71.02 5.74 15.93 0.78 6.53 7.31 3.33
58.09 12.20 22.28 0.27 7.16 7.43 3.14
Vicuna- 7B $\checkmark$ 75.94 5.35 14.71 0.53 3.48 4.01 3.50
72.63 9.76 11.65 0.81 5.15 5.96 3.43
Vicuna- 13B $\checkmark$ 85.86 2.62 9.69 0.26 1.57 1.83 3.70
80.62 10.59 5.94 0.78 2.07 2.85 3.67
[…]
Models Benign Guideline Jailbreak Dataset Average
DAN SAP200 JailbreakChatGPT T-Jailbreak
Llama2-7B-Chat $\checkmark$ 1.49 0.19 0.00 5.98 1.92
0.00 0.00 0.00 0.09 0.02
Vicuna-7B $\checkmark$ 14.90 2.00 3.95 6.24 6.77
12.48 0.12 10.53 2.73 6.47
Vicuna-13B $\checkmark$ 14.15 4.63 1.32 16.53 9.16
7.50 1.00 6.58 4.75 4.96
@@
The impact of the number of guidelines
Models Max Number Jailbreak Dataset Average
DAN SAP200 JailbreakChatGPT T-Jailbreak
Llama2-7B-Chat n=3 1.49 0.19 0.00 5.98 1.92
n=5 1.30 0.25 0.00 5.10 1.66
n=7 0.74 0.13 0.00 5.19 1.52
Vicuna-7B n=3 14.90 2.00 3.95 6.24 6.77
n=5 14.15 2.25 5.26 7.48 7.29
n=7 15.83 1.75 6.58 7.56 7.93
Vicuna-13B n=3 14.15 4.63 1.32 16.53 9.16
n=5 14.90 5.50 1.32 16.45 9.54
n=7 18.25 5.75 1.32 19.00 11.08
[…]
Model Max Number A B C D E FRR Score
Llama2- 7B-Chat n=3 71.02 5.74 15.93 0.78 6.53 7.31 3.33
n=5 64.21 8.16 21.84 0.00 5.79 5.79 3.25
n=7 68.56 7.22 18.81 1.03 4.38 5.41 3.34
Vicuna- 7B n=3 75.94 5.35 14.71 0.53 3.48 4.01 3.50
n=5 75.20 4.09 14.99 1.36 4.36 5.72 3.44
n=7 74.54 5.04 15.12 0.53 3.48 4.01 3.44
Vicuna- 13B n=3 85.86 2.62 9.69 0.26 1.57 1.83 3.70
n=5 87.70 3.28 7.38 0.82 0.82 1.64 3.76
n=7 86.13 3.47 8.53 0.53 1.33 1.83 3.72
Table 8: The results of the helpfulness scores for different maximum number of guidelines and Vanilla. The results for A, B, C, D, E, and FRR are shown as percentages (%).
@@
To evaluate the jailbreak capabilities of various techniques within the T-Jailbreak dataset, we select 30 questions per jailbreak technique and employ two defense methods: No Defense and Self-Reminder. The results are shown in Table 9. Due to space limitations, we only present the top ASR jailbreak techniques.
Our observations indicate that “Role Play” and “Logical Reasoning” are the most effective jailbreak techniques, with ASR exceeding 70% under the No Defense defense method. And under the Self-Reminder defense method, the ASR for Vicuna-7B and Vicuna-13B is also over 45%. We hypothesize that this might be due to the strong capabilities in instruction adherence and chain-of-thought of LLMs. These powerful abilities make LLMs more inclined to focus on the content provided by “Role Play” and “Logical Reasoning” jailbreak techniques, thereby overlooking the harmfulness of the queries.
These results suggest that certain jailbreak techniques exploit the inherent strengths of LLMs to bypass safety measures, highlighting the need for continuous advancement in defense strategies to detect and mitigate such vulnerabilities.
@@ Acknowledgments
We want to thank all the anonymous reviewers for their valuable comments. This paper was supported by Shenzhen Science and Technology Program (ZDSYS20230626091203008) and National Natural Science Foundation of China (62406188).

### v2

@@ Abstract
Despite being empowered with alignment mechanisms, large language models (LLMs) are increasingly vulnerable to emerging jailbreak attacks that can compromise their alignment mechanisms. This vulnerability poses significant risks to real-world applications. Existing work faces challenges in both training efficiency and generalization capabilities (i.e., Reinforcement Learning from Human Feedback and Red-Teaming). Developing effective strategies to enable LLMs to resist continuously evolving jailbreak attempts represents a significant challenge. To address this challenge, we propose a novel defensive paradigm called GuidelineLLM, which assists LLMs in recognizing queries that may have harmful content. Before LLMs respond to a query, GuidelineLLM first identifies potential risks associated with the query, summarizes these risks into guideline suggestions, and then feeds these guidelines to the responding LLMs. Importantly, our approach eliminates the necessity for additional safety fine-tuning of the LLMs themselves; only the GuidelineLLM requires fine-tuning. This characteristic enhances the general applicability of GuidelineLLM across various LLMs. Experimental results demonstrate that GuidelineLLM can significantly reduce the attack success rate (ASR) against LLM (an average reduction of 34.17% ASR) while maintaining the usefulness of LLM in handling benign queries.
Codehttps://github.com/sqzhang-lazy/GuidelineLLM
Datasetshttps://github.com/sqzhang-lazy/GuidelineLLM
Extended versionhttps://arxiv.org/pdf/2412.10423
@@ Introduction
To enhance the safety of LLMs, common approaches involve reinforcing their alignment mechanisms. Techniques such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022a; Jiang et al. 2024b) and Red-Teaming (Ganguli et al. 2022; Perez et al. 2022; Touvron et al. 2023) have proven effective in improving the ability of LLMs to detect harmful content. However, these methods demand extensive training data and significant computational resources, which can limit the speed of iterative improvements. Furthermore, additional retraining might degrade the performance of established safety alignment mechanisms (Qi et al. 2023; Zhou et al. 2024). Therefore, it is critical to investigate methodologies that can mitigate the production of harmful content without requiring modifications to the model parameters.
@@
Two prevalent methods are currently employed to enhance the safety policies of LLMs: alignment and red teaming. The objective of alignment is to bridge the gap between LLMs’ language modeling goals, such as predicting the next token during pre-training, and their ultimate aim of “following instructions and being helpful, truthful, and harmless” in practical applications (Ouyang et al. 2022). Common techniques in this domain include Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022a), Constitutional AI (Bai et al. 2022b), and self-alignment (Sun et al. 2024). These approaches focus on embedding alignment rules within pre-trained models to restrict harmful behaviors during inference. In LLMs research, the term “Red-Teaming” has recently come to denote the systematic testing or probing of LLMs to identify their potential for harmful behavior and uncover safety vulnerabilities. The critical aspect of red teaming is the aggregation of sufficient and diverse attacks, enabling LLMs to refuse responses to these attacks. Current efforts in this area mainly concentrate on automating the generation of high-quality, provocative queries, allowing LLMs to encounter a wider variety and richer range of attacks, thereby bolstering their defensive capabilities (Deng et al. 2023a; Ge et al. 2023; Jiang et al. 2024a; Zhang et al. 2024a; Xu et al. 2024). However, recent studies also demonstrate that retraining the model can lead to the deterioration of safety alignment mechanisms (Qi et al. 2023; Zhou et al. 2024). This indicates that arbitrary parameter updates to the model are not advisable and can slow the model iteration process.
@@
Ensuring that LLMs output only safe content during inference is critically important. Common methodologies to achieve this include preprocessing the input before inference, employing specialized inference paradigms, and detecting harmful content through input-output result analysis. Given the sensitivity of LLMs to input prompts, modifying the prompt content can enhance their ability to focus on harmful content. For instance, incorporating a system prompt can assist LLMs in identifying harmful content (Xie et al. 2023). By leveraging LLMs’ in-context learning capabilities, models can be conditioned to recognize and refuse harmful content (Wei, Wang, and Wang 2023). Additionally, creating multiple perturbed samples of the input and aggregating the results has proven effective (Robey et al. 2023; Cao et al. 2023). Some approaches involve having LLMs analyze the intent of the question before providing an answer (Zhang et al. 2024b; Zeng et al. 2024b). Alternatively, decoding strategies can be applied to suppress harmful prompts during generation, thereby ensuring safer outputs (Zhong et al. 2024). In terms of detection and classification, Meta has specifically fine-tuned a model named Llama-Guard, which can classify inputs and input-output pairs according to taxonomy and risk guidelines (Inan et al. 2023). In this work, we primarily focus on safeguarding LLMs during the inference phase. We aim to find a defensive paradigm that applies to different LLMs and can be continuously expanded to counter the ever-evolving jailbreak attacks.
@@
The prompts for fine-tuning GuidelineLLM include jailbreak and benign queries as inputs, with corresponding guidelines as outputs. For JailbreakLLM, prompts require generating attack queries and jailbreak queries based on specific techniques. To enhance data diversity, two examples of the same jailbreak technique are sampled as inputs during fine-tuning. During inference, both GuidelineLLM and JailbreakLLM use prompts consistent with those in the fine-tuning phase.
@@
For harmful datasets, we select three in-the-wild jailbreak datasets (i.e., DAN (Shen et al. 2023), SAP200 (Deng et al. 2023a), JailbreakChatGPT (Liu et al. 2023b)) and a popular optimization-based automatic jailbreak method (i.e., AutoDAN (Liu et al. 2023a)). Additionally, we use gpt-3.5-turbo-0125 to generate a T-jailbreak dataset comprising 1,137 queries. Besides, to evaluate the impact of GuidelineLLM on the helpfulness, we sample 400 queries from the Alpaca and GPTTeacher datasets (Taori et al. 2023).
@@
Llama2- 7B-Chat T-Jailbreak 0 1.73 5.17 0 93.1 0.16
@@
Baseline
We select Llama2-7B-Chat (Touvron et al. 2023), Vicuna-7B, and Vicuna-13B (Chiang et al. 2023) as our responding LLMs. We compare the performance of our GuidelineLLM method with several other approaches: No defense, Self-Reminderr (Xie et al. 2023), IA (Zhang et al. 2024b), and Llama3-Guard (Inan et al. 2023). For the Llama3-Guard, we only evaluate the harmfulness of the inputs. Considering that IA’s defensive method requires two rounds of inference, and both Self-Reminder and GuidelineLLM add additional prompts to the original query (i.e., Self-Reminder adds a system prompt, and GuidelineLLM adds guidelines), in the AutoDAN experiment, we use the input content from the last round of inference for each defensive method as the attack target.
Main Results
Model Method A B C D E FRR Score
Llama2 7B-Chat Va 91.62 3.14 3.4 0.26 1.57 1.83 3.82
SR 83.50 3.81 7.61 0.51 4.57 5.07 3.61
IA 31.78 14.21 16.28 2.58 35.14 37.72 2.05
GL 71.02 5.74 15.93 0.78 6.53 7.31 3.33
Vicuna 7B Va 91.27 0.79 5.03 0.26 2.65 2.91 3.77
SR 90.86 4.31 1.78 1.02 2.03 3.05 3.81
IA 77.02 6.27 6.79 0.26 9.66 9.92 3.41
GL 75.94 5.35 14.71 0.53 3.48 4.01 3.50
Vicuna 13B Va 93.62 1.33 3.72 0.53 0.8 1.33 3.86
SR 93.42 4.30 1.01 0 1.27 1.27 3.89
IA 69.67 5.19 8.47 0.82 15.85 16.67 3.12
GL 85.86 2.62 9.69 0.26 1.57 1.83 3.70
Table 4: The results of the helpfulness scores for Vanilla(Va), Self-Reminder(SR), IA and GuidelineLLM(GL). The results for A, B, C, D, E, and FRR are shown as percentages (%).
Models Base Model Jailbreak Dataset Average Helpfulness
DAN SAP200 JailbreakChatGPT T-Jailbreak FRR Score
Llama2 -7B-Chat Qwen2.5-1.5B 1.12 0.38 0.00 5.46 1.74 3.41 3.34
Qwen2.5-3B 1.73 0.31 0.00 6.16 2.02 7.42 3.25
Llama2-7B-Chat 1.49 0.19 0.00 5.98 1.92 7.31 3.33
Vicuna- 7B Qwen2.5-1.5B 19.59 2.25 9.21 10.83 10.47 3.98 3.56
Qwen2.5-3B 22.65 4.19 5.41 10.47 10.68 3.41 3.50
Llama2-7B-Chat 14.90 2.00 3.95 6.24 6.77 4.01 3.50
Vicuna- 13B Qwen2.5-1.5B 18.47 8.63 3.95 27.46 14.63 2.05 3.77
Qwen2.5-3B 21.88 7.37 5.41 28.94 15.9 1.3 3.72
Llama2-7B-Chat 14.15 4.63 1.32 16.53 9.16 1.83 3.70
[…]
@@
Models Benign Jailbreak Dataset Average Helpfulness
Guideline DAN SAP200 JailbreakChatGPT T-Jailbreak FRR Score
Llama2-7B-Chat $\checkmark$ 1.49 0.19 0.00 5.98 1.92 7.31 3.33
0.00 0.00 0.00 0.09 0.02 7.43 3.14
Vicuna-7B $\checkmark$ 14.90 2.00 3.95 6.24 6.77 4.01 3.50
12.48 0.12 10.53 2.73 6.47 5.96 3.43
Vicuna-13B $\checkmark$ 14.15 4.63 1.32 16.53 9.16 1.83 3.70
7.50 1.00 6.58 4.75 4.96 2.85 3.67
@@
The impact of base model parameters on GuidelineLLM
To explore the impact of the base model on GuidelineLLM, we conduct additional experiments by using Lora to finetune Qwen2.5-1.5B-Instruct and Qwen2.5-3B-Instruct as our GuidelineLLM. The results are shown in Table  5, and we have the following two findings:
(i) Models with fewer parameters are equally effective when trained as GuidelineLLM, and the performance of Qwen2.5-1.5B-Instruct is even better than that of Qwen2.5-3B-Instruct. This indicates that training an effective GuidelineLLM does not require a large number of parameters in the base model, which is very advantageous for deployment in practical applications.
(ii) We train two smaller LLMs using the training dataset from the main results, and found that the instruction-following ability do not achieve the best results. This suggests that smaller LLMs require a larger training dataset or full-parameter training for better performance. This explains the performance decline in Qwen2.5-1.5B-Instruct and Qwen2.5-3B-Instruct.
@@
To evaluate jailbreak capabilities in the T-Jailbreak dataset, we selected 30 questions per technique and used two defense methods: No Defense and Self-Reminder. Results (Table 9) show that ”Role Play” and ”Logical Reasoning” are the most effective techniques, likely due to LLMs’ strong instruction adherence and chain-of-thought capabilities, which may cause them to overlook harmful queries. These findings highlight how certain techniques exploit LLMs’ strengths, emphasizing the need for improved defense strategies to address such vulnerabilities.
The impact of the number of guidelines
Models Max Jailbreak Dataset Average Helpfulness
DAN SAP200 JailbreakChatGPT T-Jailbreak FRR Score
Llama2-7B-Chat n=3 1.49 0.19 0.00 5.98 1.92 7.31 3.33
n=5 1.30 0.25 0.00 5.10 1.66 5.79 3.25
n=7 0.74 0.13 0.00 5.19 1.52 5.41 3.34
Vicuna-7B n=3 14.90 2.00 3.95 6.24 6.77 4.01 3.50
n=5 14.15 2.25 5.26 7.48 7.29 5.72 3.44
n=7 15.83 1.75 6.58 7.56 7.93 4.01 3.44
Vicuna-13B n=3 14.15 4.63 1.32 16.53 9.16 1.83 3.70
n=5 14.90 5.50 1.32 16.45 9.54 1.64 3.76
n=7 18.25 5.75 1.32 19.00 11.08 1.83 3.72
[…]
@@ Acknowledgments
We want to thank all the anonymous reviewers for their valuable comments. This work was supported by National Natural Science Foundation of China (U23B2055, 62376075, 62276077, and 62406188).